OpenStudy (volleyballlover55):

The legs of a right triangle are 3 units and 6 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. 5.20 units 6.71 units 9.00 units 18.00 units

OpenStudy (volleyballlover55):

@Jadeishere

OpenStudy (mayankdevnani):



OpenStudy (mayankdevnani):

apply pythagoras theorem

OpenStudy (volleyballlover55):

i got b!

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

you're correct

OpenStudy (volleyballlover55):

ok awesome! can you please help me with more!

OpenStudy (mayankdevnani):

sure

OpenStudy (volleyballlover55):

ok great

OpenStudy (volleyballlover55):

What is the volume of a cone (in cubic inches) with a radius of 6 inches and a height of 3 inches? Use 3.14 for π. Round your answer to the nearest hundredth. 37.68 cubic inches 56.52 cubic inches 100.48 cubic inches 113.04 cubic inches

OpenStudy (mayankdevnani):

do you know the formula of volume of cone ?

OpenStudy (volleyballlover55):

yes..

OpenStudy (mayankdevnani):

what is it?

OpenStudy (volleyballlover55):

v=pi*r^2 h/3

OpenStudy (mayankdevnani):

plug in the values

OpenStudy (volleyballlover55):

how do i do that

OpenStudy (mayankdevnani):

radius=r=6 inches height=h=3 inches

OpenStudy (mayankdevnani):

and pi=3.14

OpenStudy (mayankdevnani):

\[\large \bf Volume= \frac{3.14 \times (6)^2 \times 3}{3}\]

OpenStudy (volleyballlover55):

113.04!!

OpenStudy (mayankdevnani):

use distance formula to find out the distance between warehouse and airport

OpenStudy (volleyballlover55):

ok and how do i do that..

OpenStudy (mayankdevnani):

if two points are given , (x1,y1) and (x2,y2), \[\large \bf Distance=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (volleyballlover55):

ok so what do i put on x^2 - x^1?

OpenStudy (mayankdevnani):

(x1,y1)=(24,-32) and (x2,y2)=(-24,32)

OpenStudy (volleyballlover55):

so what do i solve again?

OpenStudy (mayankdevnani):

plug in the values

OpenStudy (mayankdevnani):

\[\large \bf Distance=\sqrt{(-24-24)^2+(32+32)^2}\]

OpenStudy (volleyballlover55):

oh ok!! :)

OpenStudy (mayankdevnani):

first solve it then i'll tell what to do next

OpenStudy (volleyballlover55):

ok!

OpenStudy (volleyballlover55):

i got 80! :)

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

okay

OpenStudy (mayankdevnani):

now it travelled to factory

OpenStudy (mayankdevnani):

which is 64 miles due north of warehouse

OpenStudy (mayankdevnani):

can you guess its co-ordinates ?

OpenStudy (volleyballlover55):

i dont know... :/

OpenStudy (mayankdevnani):

add 62 miles due north it's simple

OpenStudy (volleyballlover55):

so i do 64+62?

OpenStudy (mayankdevnani):

nope

OpenStudy (mayankdevnani):

let me tell you

OpenStudy (volleyballlover55):

ok

OpenStudy (mayankdevnani):